Find the inverse of matrix A given below using the formula A -1 =CT/│A│

A="\begin{vmatrix}
 -6 & -4 & -4 \\
 -1 & -4 & 6 \\\
-6 & 5 & -1
\end{vmatrix}"


1
Expert's answer
2022-03-19T02:39:34-0400

At first step we compute the determinant of "A": "|A|=(-6)(-4)(-1)+(-1)(-4)5+(-4)(-6)6-(-6)\,6\,5+(-1)(-4)+6(-4)(-4)=-24+20+144+180+4+96=420"

Matrix "C^{\top}=\left(\begin{array}{lll}-26&-24&-40\\-37&-18&40\\-{29}&54&20\end{array}\right)"has the form: Matrix "A^{\top}" has the form: "A=\left(\begin{array}{lll}-\frac{13}{210}&-\frac{2}{35}&-\frac{2}{21}\\-\frac{37}{420}&-\frac{3}{70}&\frac{2}{21}\\-\frac{29}{420}&\frac{9}{70}&\frac{1}{21}\end{array}\right)".